OpenStudy (anonymous):

is 390 divisible by 2 3 5 6 9 10

OpenStudy (anonymous):

It is not divisible by 9.

OpenStudy (anonymous):

All others are factors.

OpenStudy (anonymous):

what about the others

OpenStudy (anonymous):

All others work

OpenStudy (anonymous):

All others work.

OpenStudy (anonymous):

what do you mean they work

OpenStudy (anonymous):

390/10=39, 390/9=43.33, 390/5=78, 390/6=65, 390/3=130.

OpenStudy (anonymous):

It means 390 is divisible by all others exept for 9.

OpenStudy (anonymous):

and I forgot, 390/2=195.

OpenStudy (anonymous):

just say say or no is the rest divisible by 390 or not